Need as much details as possible. Microeconomics.

A competitive industry consists of identical firms. Each firm has the long run total cost function TC(q)=18+½q2. If the market demand is Q(p)= 420 - p, what is the equilibrium quantity produced by each firm in the long run?

a. 12

b. 18

c. 9

d. 6
Answer #1

Answer:

Given

TC(q)=18+½q2

Q(p)= 420 - p

MC = q

equilibrium quantity produced by each firm in the long run is:

The long-run equilibrium price is that price that results in the representative firm earning zero economic profit. This will occur when MC = ATC for the representative firm.

ATC = TC / q

= 18+½q2  / q

= 18/q + q/2q

MC = q

=> 18/q + q2/2q = q

=> 18/q + q/2 = q

=> (36 +q2) = 2q2

=> 36 = 2q2 - q2

=> q = 6

Hence ,

the equilibrium quantity produced by each firm in the long run is 6

option (D) : 6 is correct.